### v1

# Spectral Differential Network Analysis for High-Dimensional Time Series

Abstract
Spectral networks derived from multivariate time series data arise in many domains, from brain science to Earth science. Often, it is of interest to study how these networks change under different conditions. For instance, to better understand epilepsy, it would be interesting to capture the changes in the brain connectivity network as a patient experiences a seizure, using electroencephalography data. A common approach relies on estimating the networks in each condition and calculating their difference. Such estimates may behave poorly in high dimensions as the networks themselves may not be sparse in structure while their difference may be. We build upon this observation to develop an estimator of the difference in inverse spectral densities across two conditions. Using an $\ell_{1}$ penalty on the difference, consistency is established by only requiring the difference to be sparse. We illustrate the method on synthetic data experiments, on experiments with electroencephalography data, and on experiments with optogentic stimulation and micro-electrocorticography data.
1 Introduction
Spectral network analysis of multivariate time series plays a key role in fields ranging from oceanography and seismology to neuroscience (Laurindo et al., 2019; James et al., 2017; Bloch et al., 2022). Studying individual networks can provide insights into how features interact; however, it is often of interest to study how networks change across conditions or in response to an external intervention. In neuroscience, for example, many neurodegenerative disorders are associated with abnormal brain connectivity networks (Bloch et al., 2022). Spectral features are regularly used to interpret many types of neuroscientific data, from electroencephalography to magnetoencephalography data (Gnassounou et al., 2023; Richard et al., 2020; Dupré la Tour et al., 2018).
Coherence, the frequency domain analog to correlation, is a common choice for investigating interactions in multivariate time series analysis. This notion is especially appealing in neuroscience applications, where activities captured at different frequencies better reveal brain oscillations in sensory-cognitive processes. Therefore, despite the availability of nonlinear association measures, such as mutual information (Belghazi et al., 2018) and transfer entropy (Ursino et al., 2020), coherence is commonly used by neuroscientists to define brain functional connectivity networks.
Similar to correlation, coherence includes the indirect effects of other nodes in the network. Thus, coherence may not be an informative measure of the dependence between nodes. A more direct measure is the inverse spectral density. The inverse spectral density between nodes $i$ and $j$ is a rescaling of the coherence after removing the linear effects of all other nodes $\{k\neq i,j\}$ (Dahlhaus, 2000). Therefore, the inverse spectral density better resembles the effective connectivity (Friston, 2011) between brain regions, providing an initial understanding of how two regions may be causally related.
Modern data collection methods have facilitated the collection of data where the dimensionality is much greater than the number of data points. This setting is often referred to as high-dimensional ($p\gg n$). When analyzing high-dimensional time series data, regularization techniques, such as the LASSO, are needed to make the problem computationally and statistically tractable (Banerjee et al., 2008). Similarly, in the frequency domain, regularization can allow for better numerical stability and better performance (Böhm and von Sachs, 2009).
While several methods exist to directly estimate the difference in inverse covariance matrices (see Tsai et al. (2022) and references therein), no such methods exist for the difference in inverse spectral densities. Naïvely, to estimate the difference in inverse spectral densities, one could first estimate the inverse spectral density in each condition and then take their difference. In high-dimensions, consistency of the resulting estimate requires both networks to be sparse (Deb et al., 2024). This assumption may be unrealistic due to the presence of hub nodes (Wang et al., 2021) and thereby lead to degraded statistical performance.
In this paper, we propose Spectral D-trace Difference (SDD), a direct estimator of the difference in inverse spectral densities, which, to our knowledge, is the first of its kind. Our direct estimator only requires sparsity of the difference (Yuan et al., 2017), a more realistic assumption. For instance, despite identifying many nonzero coherence values in resting state and stimulation states, Bloch et al. (2022) found many coherence changes to be near zero. These small differences are likely due to noise in the estimation procedure, corresponding to no underlying change in coherence. Assuming sparsity of the difference, and only under a geometrically decaying time dependence condition, we show that our direct difference estimator consistently estimates the true difference. Performance of our estimator is studied in a variety of simulation settings, as well as two real data applications.
Notation.
We use $\|A\|_{1}$, $\|A\|_{F}$, $\|A\|_{\infty}$, $\|A\|_{0}$ to denote the $\ell_{1}$, Frobenius, infinity, and $\ell_{0}$ norms of a matrix $A$, respectively. The element-wise absolute value of a matrix $A$ is denoted as $|A|$. The conjugate transpose of a matrix or vector will be denoted as $A^{H}$. We denote the inner product between two matrices $A$ and $B$ as $\langle A,B\rangle=\mathrm{Tr}(AB^{T})$ and use $A\otimes B$ to represent their Kronecker product.
2 Method
Figure 1 provides a visual explanation of our method, SDD, and each step (A)–(E) is elaborated upon in this section.
Figure 1: SDD Method. (A) Inputs: Time series datasets are observed for two conditions. (B) Computing the Spectral Densities: We then compute the spectral densities in each condition using the smoothed periodogram. (C) Expanding to Real Space: These spectral densities are expanded to the real space. (D) Direct Estimation of the Difference in Inverse Spectral Densities: Using the expanded spectral densities we directly compute the difference estimator. (E) Output: We form the difference in inverse spectral densitites by taking the (1,1) and (2,1) blocks of $\hat{\Delta}$.
(A) Inputs.
Suppose we have mean-zero stationary data in two conditions, denoted as 1 and 2. In condition 1, there are $n_{1}$ observations of $p$ covariates, while condition 2 has $n_{2}$ observations. For example, in neuroscience applications studying how stimulation changes the brain network as in Bloch et al. (2022), conditions 1 and 2 might represent pre- and post-stimulation, respectively.
(B) Computing the Spectral Densities.
The spectral density in condition $l\in\{1,2\}$ at frequency $\lambda$, $f_{l}(\lambda)$, is defined as the Fourier transform of the autocovariance matrix. Intuitively, the spectral density represents the portion of the covariance of the signal that can be attributed to the specific frequency $\lambda$ (Brockwell and Davis, 1991, pp. 331–332). In practice, many different $\lambda$ values are of interest, but for notational simplicity we assume $\lambda$ is fixed throughout and suppress the dependence on $\lambda$. Mathematically we can write our data as $\{\mathbf{x}_{l,t}\}$ for $t=1,\dots,n_{l}$ where $\mathbf{x}_{l,t}$ is a $p$-dimensional vector in condition $l\in\{1,2\}$. For each condition $l$, the autocovariance at lag $h\in\mathbb{Z}$ is denoted as $\Gamma_{l}(h)=\mathbb{E}\left(\mathbf{x}_{l,t}\mathbf{x}^{T}_{l,t+h}\right)$. If $\sum_{h=-\infty}^{\infty}\left|\Gamma_{l}(h)\right|<\infty$ then the spectral density in condition $l$ at frequency $\lambda$ exists and is defined as
$f_{l}(\lambda)=\frac{1}{2\pi}\sum_{h=-\infty}^{\infty}e^{-i\lambda h}\Gamma_{l}(h),\qquad-\pi\leq\lambda\leq\pi.$
To estimate the difference of inverse spectral densities, we will require estimates of the spectral densities themselves. We use the smoothed periodogram to estimate the spectral density matrix and refer to it as $\hat{f}_{l}$. To define the smoothed periodogram, we must first define the periodogram. For condition $l$, the periodogram at Fourier frequency $\{\lambda_{j}=2\pi j/n_{l},-\lfloor(n_{l}-1)/2\rfloor\leq j\leq\lfloor n_{l}/2\rfloor\}$ is defined as
$P_{l}(\lambda_{j})=\frac{\left(\sum_{t=1}^{n_{l}}\mathbf{x}_{l,t}e^{-i\lambda_{j}t}\right)\left(\sum_{t=1}^{n_{l}}\mathbf{x}_{l,t}e^{-i\lambda_{j}t}\right)^{H}}{2\pi n_{l}}\,.$
Recall $\mathbf{x}^{H}$ is the conjugate transpose of the vector $\mathbf{x}$. Using the Fast Fourier transform (FFT) algorithm, the periodogram can be computed quickly. The periodogram is known to be an inconsistent estimator of the spectral density so a smoothed version is used where the periodograms of $2M_{n_{l}}$ nearby frequencies are averaged (Brockwell and Davis, 1991, pp. 347–350). $M_{n_{l}}$ is referred to as the bandwidth or smoothing span. That is, the smoothed periodogram at fourier frequency $\lambda_{j}$ is generated as
$\hat{f}_{l}(\lambda_{j})=\frac{1}{2M_{n_{l}}+1}\sum_{k=j-M_{n_{l}}}^{j+M_{n_{l}}}P_{l}(\lambda_{k}).$
(C) Expanding to the Real Space.
The spectral density, $f_{l}$, is complex-valued in general which can complicate the estimation. To transform this problem to the real space, we note that any complex matrix can be written as the sum of a real matrix and $i$ times another real matrix (for $i^{2}=-1$). Specifically we write the spectral density and its inverse as $f_{l}=A_{l}+iB_{l}$, and $f_{l}^{-1}=\tilde{A}_{l}+i\tilde{B}_{l}$ respectively where $A$, $B$, $\tilde{A}$, and $\tilde{B}$ are real matrices.
With these representations, we can now work in the real space by studying $\Sigma_{l}=\begin{bmatrix}A_{l}&-B_{l}\\ B_{l}&A_{l}\end{bmatrix}\in\mathbb{R}^{2p\times 2p}$ instead of $f_{l}$. Then, by Lemma A.1 of Fiecas et al. (2019),
$\begin{bmatrix}A_{l}&-B_{l}\\ B_{l}&A_{l}\end{bmatrix}\begin{bmatrix}\tilde{A}_{l}&-\tilde{B}_{l}\\ \tilde{B}_{l}&\tilde{A}_{l}\end{bmatrix}=\Sigma_{l}\Sigma_{l}^{-1}=I_{2p},$
where $I_{2p}$ represents the $2p\times 2p$ identity matrix. It is worth noting that $\Sigma_{l}^{-1}$ is the expansion of $f_{l}^{-1}$ to the real space. Moreover, $f_{l}$ and $f_{l}^{-1}$ can easily be recovered by taking the (1,1) and (2,1) blocks of $\Sigma_{l}$ and $\Sigma_{l}^{-1}$, respectively.
(D) Direct Estimation of the Difference in Inverse Spectral Densities.
Having defined the problem in the real space, we use the $\ell_{1}$ penalized D-trace loss function (Yuan et al., 2017) to directly estimate $\Delta=\Sigma_{1}^{-1}-\Sigma_{2}^{-1}$.
To justify this estimation strategy, in Appendix A.1.1 we show that this loss is convex and the population loss is minimized at the true $\Delta=\Sigma_{1}^{-1}-\Sigma_{2}^{-1}$. The inputs to this estimator are $\hat{\Sigma}_{1}$ and $\hat{\Sigma}_{2}$ which are formed by expanding $\hat{f}_{1}$ and $\hat{f}_{2}$ to the real space respectively. Our estimator is
$\hat{\Delta}=\operatorname*{arg\,min}_{\Delta}\text{\hskip 3.55658pt}\frac{1}{4}\left(\langle\hat{\Sigma}_{2}\Delta,\Delta\hat{\Sigma}_{1}\rangle+\langle\hat{\Sigma}_{1}\Delta,\Delta,\hat{\Sigma}_{2}\rangle\right)-\langle\Delta,\hat{\Sigma}_{2}-\hat{\Sigma}_{1}\rangle+\tau_{n_{1},n_{2}}\|\Delta\|_{1}$ (1)
where $\tau_{n_{1},n_{2}}$ is a penalty parameter depending on sample sizes $n_{1},n_{2}$ and $\|\Delta\|_{1}$ is the $\ell_{1}$ penalty which returns the sum of the absolute values of all entries of a matrix. We solve for $\hat{\Delta}$ using an alternating direction method of multipliers (ADMM) algorithm similar to Yuan et al. (2017).
(E) Output.
Recall that the inverse spectral density, $f_{l}^{-1}$, represents the coherence between two nodes after removing the linear effects of all other nodes. Our goal is to estimate how this effective connectivity differs between the two conditions. That is, we wish to estimate $f_{1}^{-1}-f_{2}^{-1}$. Using our representation in the real space, the true difference $f_{1}^{-1}-f_{2}^{-1}$ is recovered by taking the (1,1) and (2,1) blocks of
$\Sigma_{1}^{-1}-\Sigma_{2}^{-1}=\begin{bmatrix}\tilde{A}_{1}&-\tilde{B}_{1}\\ \tilde{B}_{1}&\tilde{A}_{1}\end{bmatrix}-\begin{bmatrix}\tilde{A}_{2}&-\tilde{B}_{2}\\ \tilde{B}_{2}&\tilde{A}_{2}\end{bmatrix}:=\Delta.$
Similarly, using our estimate $\hat{\Delta}$ of $\Delta$ we can estimate $f_{1}^{-1}-f_{2}^{-1}$ using the (1,1) and (2,1) blocks of $\hat{\Delta}$.
In the following section, we study the consistency of our estimator. Intuitively, if $\hat{f}_{l}$ estimates the spectral density well, we should be able to estimate $\Delta$ well. Our theoretical results show that this is indeed the case and the rate at which the smoothed periodogram $\hat{f}_{l}$ converges to the true periodogram is also the rate at which $\hat{\Delta}$ converges to $\Delta$. Due to the dependence in the data, the smoothed periodogram and $\hat{\Delta}$ are not $\sqrt{n}$ consistent as one might expect in the i.i.d case. Rather, we show that for an optimal choice of smoothing span, $\hat{f}_{l}$ and $\hat{\Delta}$ converge to $f_{l}$ and $\Delta$ respectively at a rate of $n^{-1/3}$.
3 Theory
In this section, we study the theoretical properties of our SDD estimator, $\hat{\Delta}$. Specifically, we establish the convergence rate of $\hat{\Delta}$ to the population quantity $\Delta$. It is easily seen in Equation 1 that our estimator $\hat{\Delta}$ relies on expanded estimates of the spectral density $\hat{\Sigma}_{1},\hat{\Sigma}_{2}$. Thus, it is natural that the rate of convergence of $\hat{\Delta}$ to $\Delta$ relies on the rates of convergence of $\hat{f}_{1}$ and $\hat{f}_{2}$ to $f_{1}$ and $f_{2}$.
Theorem 1.
Suppose Assumption 2 in Appendix A.2 holds and that $\tau_{n_{1},n_{2}}\geq O(\min(n_{1},n_{2})^{-1/3})$ and $M_{n_{l}}=O(n_{l}^{2/3})$ for $n_{1}$ and $n_{2}$ large enough. Then, for any $H>0$, there exists some $C>0$ such that with probability greater than $1-Cp^{2}\min(n_{1},n_{2})^{-H}$,
$\|\hat{\Delta}-\Delta\|_{F}\leq s_{\Delta}O\left(\min(n_{1},n_{2})^{-1/3}\right),$
where $s_{\Delta}$ is the number of non-zero elements in the true difference matrix $\Delta$.
To establish convergence rates of $\hat{f}_{1}$ and $\hat{f}_{2}$ in high-dimensions, we rely on Theorem 3.1 of Fiecas et al. (2019), which uses the functional dependency framework of Wu (2005) and Wu and Zaffaroni (2018). Specifically, our convergence rates hinge on the assumption of a geometrically decaying dependence in the data. An overview of the functional dependence framework is given in Appendix A.2 and a formal statement of the geometric decay assumption is given in Assumption 2 in Appendix A.2. The geometric decay assumption is mild and is satisfied by many processes including ARMA, ARMA-ARCH, and ARMA-GARCH processes, as well as other nonlinear autoregressive processes (Shao and Wu, 2007; Liu and Wu, 2010). A sketch of the proof for Theorem 1 is provided below while the full proof is given in Appendix A.3.
Proof sketch.
The concentration bound in Theorem 3.1 of Fiecas et al. (2019) is essential to our theoretical analysis and depends on the smoothing span $M_{n_{l}}$. We therefore begin by selecting $M_{n_{l}}$ to optimize this bound which gives $M_{n_{l}}=O(n_{l}^{2/3})$. We can then plug in this $M_{n_{l}}$ to get a version of Theorem 3.1 that only depends on constants and $n_{l}$. We let $\Lambda_{l}=\{\lambda_{j}=2\pi j/n_{l},\lfloor(n_{l}-1)/2\rfloor\leq j\leq\lfloor n_{l}/2\rfloor\}$. For each condition $l$, Theorem 3.1 of Fiecas et al. (2019) becomes, for any $H>0$ there exists a $C>0$,
$\mathbb{P}\left(\sup_{\lambda_{j}\in\Lambda_{l}}\|\hat{f}_{l}(\lambda_{j})-f_{l}(\lambda_{j})\|_{\infty}>O\big{(}n_{l}^{-1/3}\big{)}\right)\leq Cp^{2}n_{l}^{-H}\,$ (2)
where $\|A\|_{\infty}$ is the infinity norm of $A$. We proceed by using the high-dimensional M-estimator framework of Negahban et al. (2012) and similar strategies to Wang et al. (2021). In this framework, the D-trace loss in Equation 1 is our convex differentiable loss function and the $\ell_{1}$ norm is our regularization penalty. Equipped with our new concentration bound in (2), we can now prove everything needed for Corollary 1 of Negahban et al. (2012). We briefly restate some notation used below. $\|A\|_{1}$ and $\|A\|_{\infty}$ denote the $\ell_{1}$ and infinity norms of a matrix $A$ respectively while $A\otimes B$ represents the Kronecker product between $A$ and $B$. We also use $\|\mathbf{x}\|_{1}$ and $\|\mathbf{x}\|_{2}$ to denote the $\ell_{1}$ and $\ell_{2}$ norms of a vector $\mathbf{x}$.
First, we must show the restricted strong convexity (RSC) of the D-trace loss. This amounts to showing that $\mathbf{m}^{T}\left(\nabla^{2}L(\Delta,\hat{\Sigma}_{1},\hat{\Sigma}_{2})\right)\mathbf{m}>\kappa_{L}\|\mathbf{m}\|_{2}^{2}$ for all $\mathbf{m}\in\mathcal{C}(S_{\Delta})$, where $\mathcal{C}(S_{\Delta})=\left\{\mathbf{m}\in\mathbb{R}^{4p^{2}}\big{|}\|\mathbf{m}_{S_{\Delta}^{c}}\|_{1}\leq 3\|\mathbf{m}_{S_{\Delta}}\|_{1}\right\}$ for some $\kappa_{L}$. Recall that $\nabla^{2}L(\Delta,\hat{\Sigma}_{1},\hat{\Sigma}_{2})=0.5(\hat{\Sigma}_{1}\otimes\hat{\Sigma}_{2}+\hat{\Sigma_{2}}\otimes\hat{\Sigma}_{1})$. Thus, we study $\mathbf{m}^{T}(\hat{\Sigma}_{1}\otimes\hat{\Sigma}_{2})\mathbf{m}$. The same results hold for $\mathbf{m}^{T}(\hat{\Sigma}_{2}\otimes\hat{\Sigma}_{1})\mathbf{m}$. We hence get
$\mathbf{m}^{T}(\hat{\Sigma}_{1}\otimes\hat{\Sigma}_{2})\mathbf{m}\geq$ $\mathbf{m}^{T}(\Sigma_{1}\otimes\Sigma_{2})\mathbf{m}+\mathbf{m}^{T}\left(\hat{\Sigma}_{1}\otimes\hat{\Sigma}_{2}-\Sigma_{1}\otimes\Sigma_{2}\right)\mathbf{m}$
$\geq$ $\lambda_{\mathrm{min}}(\Sigma_{1})\lambda_{\mathrm{min}}(\Sigma_{2})\|\mathbf{m}\|_{2}^{2}-16s_{\Delta}\|\hat{\Sigma}_{1}\otimes\hat{\Sigma}_{2}-\Sigma_{1}\otimes\Sigma_{2}\|_{\infty}\|\mathbf{m}\|_{2}^{2},$
where $\lambda_{\mathrm{min}}(f_{l})$ is the minimum eigenvalue of $f_{l}$. For large enough $n_{l}$, $\|\hat{f}_{l}-f_{l}\|_{\infty}$ and hence $\|\hat{\Sigma}_{1}\otimes\hat{\Sigma}_{2}-\Sigma_{1}\otimes\Sigma_{2}\|_{\infty}$ are small by (2) with high probability. Thus for large enough $n_{1}$ and $n_{2}$, with high probability, we can ensure that $\mathbf{m}^{T}(\hat{\Sigma}_{1}\otimes\hat{\Sigma}_{2})\mathbf{m}\geq\frac{\lambda_{\mathrm{min}}(f_{1})\lambda_{\mathrm{min}}(f_{2})}{2}\|\mathbf{m}\|_{2}^{2}$ showing that RSC holds with $\kappa_{L}=\frac{\lambda_{\mathrm{min}}(f_{1})\lambda_{\mathrm{min}}(f_{2})}{2}$.
Next, we choose the penalty scale $\tau_{n_{1},n_{2}}$. Following Negahban et al. (2012), we require that
$\tau_{n_{1},n_{2}}$ $\geq 2\|\nabla L_{D}(\Delta,\hat{\Sigma}_{1},\hat{\Sigma}_{2})\|_{\infty}$
$=2\|0.5\left(\hat{\Sigma}_{1}\Delta\hat{\Sigma}_{2}+\hat{\Sigma}_{2}\Delta\hat{\Sigma}_{1}\right)-\left(\hat{\Sigma}_{1}-\hat{\Sigma}_{2}\right)\|_{\infty}\,.$
Let $\Gamma=0.5\left(\Sigma_{2}\otimes\Sigma_{1}+\Sigma_{1}\otimes\Sigma_{2}\right)$ and $\hat{\Gamma}=0.5\left(\hat{\Sigma}_{2}\otimes\hat{\Sigma}_{1}+\hat{\Sigma}_{1}\otimes\hat{\Sigma}_{2}\right)$. Then, we can show that
$\left\|0.5\left(\hat{\Sigma}_{1}\Delta\hat{\Sigma}_{2}+\hat{\Sigma}_{2}\Delta\hat{\Sigma}_{1}\right)-\left(\hat{\Sigma}_{1}-\hat{\Sigma}_{2}\right)\right\|_{\infty}\leq\left\|\hat{\Gamma}-\Gamma\right\|_{\infty}\left\|\Delta\right\|_{1}+\left\|\hat{\Sigma}_{1}-\Sigma_{1}\right\|_{\infty}+\left\|\hat{\Sigma}_{2}-\Sigma_{2}\right\|_{\infty}$
Similar to RSC, we can show that, for large enough $n_{1}$ and $n_{2}$, with high probability,
$\left\|0.5\left(\hat{\Sigma}_{1}\Delta\hat{\Sigma}_{2}+\hat{\Sigma}_{2}\Delta\hat{\Sigma}_{1}\right)-\left(\hat{\Sigma}_{1}-\hat{\Sigma}_{2}\right)\right\|_{\infty}\leq O(\min(n_{1},n_{2})^{-1/3}).$ (3)
Therefore, choosing $\tau_{n_{1},n_{2}}>O(\min(n_{1},n_{2})^{-1/3})$ the penalty selection requirements are satisfied with high probability. Theorem 1 then follows from Corollary 1 of Negahban et al. (2012). ∎
We now discuss interesting intermediate results from the proof sketch. In equation (2), we show that, with high probability and a smoothing span of $M_{n_{l}}=O(n^{2/3})$, the deviations of the smoothed periodogram from the true spectral density in condition $l$ are $O(n_{l}^{-1/3})$. Furthermore, Equation 3 establishes that, with high probability, the maximum norm of the derivative of the D-trace loss is $O(n_{l}^{-1/3})$.
Due to the dependence in the data, the convergence rate in Theorem 1 is slower than the $n^{-1/2}$ rates for i.i.d data. In fact, as we choose $M_{n}$ to minimize the deviations in Theorem 3.1 of Fiecas et al. (2019), our rate is the best we can expect to achieve using this general dependence framework when only assuming a geometrically decaying time dependence.
4 Simulations
In this section, we evaluate the performance of the proposed estimator $\hat{\Delta}$ using simulation studies. All simulations use VAR$(1)$ processes, with $p=54$, as this allows us to compute the true spectral density using results from Sun et al. (2018). We control the sparsity of the difference in inverse spectral densities by using a block diagonal structure. In all settings, the transition matrix in condition 1 is the same as that in condition 2 except the last $3\times 3$ block is multiplied by $-1$. The first setting uses similar coefficients as in Sun et al. (2018). In the second and third simulation settings, the transition matrix consists of one $51\times 51$ block and one $3\times 3$ block. The larger block was generated with 60% and 95% sparsity in the second and third settings, respectively. All simulations were performed using $n=100,200,500,1000,2000$ observations for both conditions. Smoothed periodograms were computed from data for each condition using a smoothing window of $M_{n}=\lceil n^{2/3}\rceil$ and then converted to the real space to generate $\hat{\Sigma}_{l}$. All computations were performed on a MacBook with a 2.3 GHz Quad-Core Intel Core i7 processor and 16 GB of memory. More information on the simulation setup, optimization techniques, and tuning parameter selection can be found in Appendix A.4.
Figure 2: Simulation 1. Results are reported as mean (dots) and SE (vertical lines) where the mean and SE are taken across all frequencies for a given sample size $n$.
As previously noted, our method is the first to directly estimate the difference in inverse spectral densities and no direct competitors or state-of-the-art methods exist. Therefore, we conceptualized two potential alternatives to our SDD method and compare all methods in our simulations. The first alternative method is the Naïve method, which estimates $\Delta$ by taking the difference after estimating individual inverse spectral densities separately, $\hat{\Delta}_{\mathrm{N}}=\hat{\Sigma}_{1}^{-1}-\hat{\Sigma}_{2}^{-1}$. For the Naïve method, the graphical LASSO (GLASSO, Friedman et al., 2008) was used to estimate a sparse inverse spectral density in each condition. Specifically we used the fast implementation from Sustik and Calderhead (2012) available in the glassoFast package in R (GPL ($\geq$ 3.0)). To induce sparsity in individual estimates and (potentially) their difference, GLASSO was tuned for each condition separately using eBIC with $\gamma=0.5$ (Foygel and Drton, 2010). The next method directly thresholds the difference to induce sparse estimates. Specifically, we estimate the inverse spectral density in each condition and take the difference and then use Hard thresholding. The thresholds were tuned via BIC. Similar to Deb et al. (2024), all methods are compared based on the following metrics: Accuracy, Precision, Recall, and Relative Root Mean Square Error (RRMSE). The full definitions can be found in Appendix A.4.
Figure 3: Simulation 2. Results are reported as mean (dots) and SE (vertical lines) where the mean and SE are taken across all frequencies for a given sample size $n$.
Simulation results for Accuracy and RRMSE in simulations 1 and 2 are reported in Figures 2 and 3. Results for Precision and Recall for simulations 1 and 2 can be found in Appendix A.4 Figures 8 and 9. Results from simulation 3 are in Appendix A.4, Figures 10 and 11. The full set of results for each of the simulations are in Appendix A.4 Tables 1, 2, 3. In general, SDD has better accuracy than the Naïve method with the difference in accuracy increasing as the sample size increases. This is likely due to the fact that, in general, SDD estimates much fewer edges than the Naïve method. This also results in fewer false positive edges but more false negatives, resulting in better precision for SDD at the expense of reduced recall; see Table 1 in Appendix A.4. Across all settings we also see that SDD generally has lower RRMSE compared to the Naïve method especially for large sample sizes. Overall, compared to the Naïve method, SDD identifies a similar number of true edges using a much smaller edge set. While Hard thresholding has slightly better accuracy than SDD in simulation setting 1 and better precision, SDD has much better RRMSE and recall across all settings and sample sizes. The standard errors of all methods are large due to the large variation across frequencies. To further support our analysis, we have also summarized the difference between SDD and each of the competing methods for each metric in Appendix A.4, Tables 456789.
5 Application to EEG Data
We apply our direct difference estimator (SSD) and competing estimators to electroencephalograms (EEG) data from Hatlestad-Hall et al. (2022). Data were recorded with a 64 channel EEG array for 111 healthy subjects at a sampling frequency of 1024 Hz. Four minutes of brain activity was recorded while subjects were resting with their eyes closed. For 42 subjects, a second session was recorded 2–3 months after the initial session. For our analysis, we used the pre-cleaned data provided in OpenNeuro Dataset ds003775 (Markiewicz et al., 2021). Specific cleaning steps can be found in Hatlestad-Hall et al. (2022). We also downsampled the data to 512 Hz.
To validate our method, we analyzed the subjects with follow-up and without follow-up separately. For those with follow-up, we estimated the difference in networks from the first to the second session (across session analysis). For those without, we estimated the difference in networks from 0-60s to 120-180s (within session analysis). The 60-120s block was used as a rest. It has been shown that brain networks are temporally dynamic (Zalesky et al., 2014; Nobukawa et al., 2019). Therefore, a priori, we expect the differential networks to be sparser in the within session analysis compared to the across session analysis and finding such results would provide validation of our method.
We compare the sparsity of the estimated differential network for SDD, the Naïve, and Hard threshold methods across the commonly-used Theta, Beta, Gamma and High-Gamma bands. Sparsity was defined as the proportion of entries in the estimated difference that were non-zero. Five or six evenly spaced frequencies were considered for each band. They were (in Hz): Theta: (4,5,6,7,8), Beta: (12, 16, 20, 24, 28), Gamma: (30, 40, 50, 60, 70), High-gamma: (80, 95, 110, 125, 140, 150). Results averaged over subjects and frequencies within each band are presented in Figures 4 and  5.
Figure 4: EEG Within Session Analysis. Vertical lines indicate SE.
Figure 5: EEG Across Session Analysis. Vertical lines indicate SE.
Figures 4 and  5 shows that SDD is the only method that estimates a sparser difference in the within session analysis compared to the across session analysis across all frequency bands. On the other hand, competing methods have similar or less sparsity in the within session analysis compared to the across-session analysis. These results clearly indicate that SDD estimates are more consistent with the experimental setting in this application.
6 Application to $\boldsymbol{\mu}\mathrm{\textbf{ECoG}}$ DATA
We next apply our SDD estimator to optogenetic stimulation micro-electrocorticography ($\mu\mathrm{ECoG}$) data from Yazdan-Shahmorad et al. (2016, 2018); Bloch et al. (2022) [^1] The data is publicly available at https://datadryad.org/stash/dataset/doi:10.7272/Q61834NF. Studying how stimulation changes the brain network using such experiments can help design protocols to change the brain network in a targeted way. This can lead to treatments for conditions, such as schizophrenia or epilepsy, which are known to stem from abnormal brain connectivity networks (Bloch et al., 2022). A series of experimental and control sessions were conducted where neural activity from the surface of the brain of two adult male rhesus macaques was recorded. There were a total of 32 experimental and 4 control sessions. Each experimental session consisted of five 10-minute stimulation blocks (stimulation state) with a 5-minute recording block (resting state) before and after. Thus, each experimental session alternated between resting state and stimulation state blocks consisting of 6 total resting state blocks and 5 total stimulation state blocks. Control sessions had the same design as the experimental sessions except no stimulation occurred. Stimulation in experimental sessions was performed using optogenetics for simultaneous stimulation and recording. For each session, two locations corresponding to two electrodes were randomly chosen for optogenetic stimulation. To stimulate the brain, a 5ms laser pulse with a delay of either 100ms or 10ms was delivered every 200ms during the stimulation block. Recordings were obtained using a $92\mathrm{mm}^{2}$ micro-electrocorticography array of 96 electrodes at 24 kHz spanning the primary motor and sensorimotor (M1 and S1) regions. We downsampled to 500 Hz for our analysis.
Prior work shows that stimulation induces changes across the entire brain connectivity network (Yazdan-Shahmorad et al., 2018; Bloch et al., 2022). We also study both how stimulation changes the connectivity between stimulation sites and across the entire network. However, instead of using a Naïve estimator of the connectivity changes as in Bloch et al. (2022) we use SDD. Using SDD, we estimate the change in inverse spectral densities from resting state to the following resting state (RS to RS) and from resting state to the following stimulation state (RS to SS). For each session there are 5 RS to RS differences (RS1 to RS2, RS2 to RS3, …, RS5 to RS6) and 5 RS to SS difference (RS1 to SS1, …, RS5 to SS5). Note that an estimated SDD coefficient of zero indicates no change while non-zero coefficients indicate a change. This could be either a change in magnitude of the connection, a new connection, or a lost connection. The frequencies we study in each band are the same as in Section 5. Results are averaged over all sessions, frequencies within each band, delay (control, 10ms, or 100ms) and RS to RS or SS to SS.
Figure 6: Probability of Change in Edges Between Stimulation Sites by Frequency Band and Delay. Vertical bars indicate standard errors.
First, we study how stimulation changes the edges between stimulation sites. We compute the probability of edge change between the stimulation sites for each band, delay, and RS to RS or RS to SS. Results are displayed in Figure 6. Compared to the control sessions, stimulation with a 10ms delay is more likely to induce a change in connectivity between the stimulation sites across all frequency bands. After stimulation ends (RS to RS analysis), this connectivity change is attenuated, but still persists. This is in line with prior work which showed a change in functional connectivity between stimulation sites in the stimulation state which persisted into the following resting state (Bloch et al., 2022). Interestingly, stimulation with a 100ms was also shown to induce a change in connectivity relative to the control sessions despite being outside the effective range for inducing a connectivity change. This is likely a result of the interaction between stimulation and the underlying network which has been shown to mediate connectivity changes (Bloch et al., 2022).
Next, we study how stimulation affects the entire brain network. Specifically we focus on how the distance of electrodes from stimulation sites impacts edge changes. While we expect electrode pairs closer to stimulation sites to have a higher likelihood of an edge change, the effect on the rest of the network is unclear as the effect of stimulation is mediated by the existing connectivity network (Bloch et al., 2022). For every electrode, we compute the distance to closest stimulation site and, for every pair of electrodes, we average these distances to obtain the average distance to the closest stimulation site. We then study how the relative probability of an edge change between electrodes varies by the mean stimulation distance. Since electrodes that are closer together are more likely to have an edge change even during control sessions (see Figure 12 in Appendix A.5), we need to adjust for this baseline probability of edge change. The relative probability of edge change thus adjusts for the baseline probability of an edge change for two electrodes at a specific distance apart. It can be interpreted as the probability of an edge change above/below what we would expect for two electrodes at the same distance apart in a control session. In Figure 7, we apply a generalized additive model (GAM) smoother to more clearly see trends in the data.
Figure 7: Relative Probability of Change in Edge by Average Distance to Stimulation Site. Lines are generated by smoothing using a GAM. Shaded areas represent 95% confidence intervals. A value of 0 (black line) represents no difference in probability of edge change relative to control sessions.
For both the 100ms and 10ms delays, we see similar curve shapes in the RS to RS and RS to SS analyses, suggesting that network changes induced during stimulation persist to the resting state. Across the Beta, Gamma, and High-Gamma bands, we see an increase in probability of an edge change for pairs of electrodes close to simulation sites for both the 10ms and 100ms delay, which is in line with prior works (Bloch et al., 2022). For medium distances, we see near zero relative probability of edge changes for all delays and all bands except Theta. However, in the Beta and Gamma bands, for pairs of electrodes that are $\approx 10-12.5$mm away from a stimulation site, the 100ms delay appears to induce more edge changes while the 10ms delay appears to have the opposite effect inducing fewer edge changes relative to baseline. In the Theta band, the 100ms delay induces fewer edge changes relative to baseline as distance from stimulation site increases. Conversely, the 10ms delay results in fewer edge changes relative to baseline in the Theta band for distances from stimulation site $<10$mm but produces more changes as distances from the stimulation site increase past 10mm. This suggests that when designing a stimulation protocol to induce connectivity changes, depending on the intended effect at different electrodes, either a 10ms or 100ms stimulation could be used.
7 Discussion
We proposed a direct estimate of the differences in inverse spectral densities between two conditions, termed the Spectral D-trace Difference (SDD) estimator. By using the direct difference estimator, we only need to assume sparsity in the difference of inverse spectral densities. Compared to the usual assumptions of sparsity of each inverse spectral density, the sparsity of the difference is more realistic if, for example, the inverse spectral densities do not change much between conditions. This is indeed what we expect in many biological settings, especially neurodegenerative disorders that are associated with changes in brain connectivity. Convergence rates of our estimator to the true difference were derived using only an assumption on decaying dependence within the time series.
The SDD estimator was compared to the Naïve method of estimating sparse inverse the spectral densities and taking their difference, as well as a method that involved estimating the inverse spectral densities, forming the difference and then Hard thresholding the difference. The three methods were compared across three different simulation settings, where the true difference in inverse spectral densities was sparse. Comparison metrics included precision, recall, accuracy, and relative root mean square error (RRMSE). When compared to the other methods, SDD performed better in accurately identifying edges and non-edges and the SDD estimates had a lower relative root mean square error.
The SDD estimator was applied to both EEG and optogenetic stimulation $\mu\mathrm{ECoG}$ data. Using the EEG data we further validated SDD by showing it estimated sparser differences between networks that were closer in time, in line with expected behavior. Lastly SDD was applied to $\mu\mathrm{ECoG}$ data with optogenetic stimulation to study how the connectivity of the brain network changes in response to stimulation. We found that different stimulation parameters affected the connectivity network differently. This difference was likely due to the fact that the underlying brain network mediates changes in response to stimulation. We also found that the effects of stimulation persisted beyond the stimulation state.
While consistency of SDD is established in Theorem 1, this result does not quantify the degree of uncertainty around these estimates. A particularly interesting future direction is developing methods for uncertainty quantification, such as confidence intervals and hypothesis testing. These may be important when the relationship between two brain regions or nodes in the graph is of primary interest or when the sample size is small and noise may be large relative to the signal.
Appendix A Appendix
[…]
1Contents
A.1 Additional Details for Methods
A.1.1 D-trace Loss
To estimate $\Delta$, we use the $\ell_{1}$ penalized D-trace loss function from Yuan et al. (2017):
$L_{D}\left(\Delta,\Sigma_{2},\Sigma_{1}\right)=\frac{1}{4}\left(\langle\Sigma_{2}\Delta,\Delta\Sigma_{1}\rangle+\langle\Sigma_{1}\Delta,\Delta\Sigma_{2}\rangle\right)-\langle\Delta,\Sigma_{2}-\Sigma_{1}\rangle\,$
where $\langle A,B\rangle=\mathrm{Tr}(AB^{T})$. Similar to Yuan et al. (2017), we can take the derivative with respect to $\Delta$ to see that $\Sigma_{1}^{-1}-\Sigma_{2}^{-1}$ minimizes the D-trace loss. More specifically, let
$\frac{\partial{L_{D}}}{\partial{\Delta}}=\left(\Sigma_{2}\Delta\Sigma_{1}+\Sigma_{1}\Delta\Sigma_{2}\right)/2-\left(\Sigma_{2}-\Sigma_{1}\right)\,.$ (4)
Evaluating $\frac{\partial{L_{D}}}{\partial{\Delta}}$ at ${\Delta=\Sigma_{1}^{-1}-\Sigma_{2}^{-1}}$ yields 00, establishing that $\Delta=\Sigma_{1}^{-1}-\Sigma_{2}^{-1}$ minimizes $L_{D}$. Furthermore, the Hessian of $L_{D}$ with respect to $\Delta$ is $\frac{\partial^{2}{L_{D}}}{\partial{\Delta^{2}}}=(\Sigma_{1}\otimes\Sigma_{2}+\Sigma_{2}\otimes\Sigma_{1})/2$ where $\otimes$ is the kronecker product. Both $\Sigma_{1}$ and $\Sigma_{2}$ are positive definite, and thus $L_{D}$ is convex in $\Delta$ and $\Delta=\Sigma_{1}^{-1}-\Sigma_{2}^{-1}$ is the unique minimizer.
In practice, the population quantites $\Sigma_{1}$ and $\Sigma_{2}$ are not available. Instead, we use the estimates $\hat{\Sigma}_{1}$ and $\hat{\Sigma}_{2}$. We can additionally incorporate an $\ell_{1}$ penalty to estimate a sparse $\Delta$ which gives us our estimator in (1).
A.2 Functional Dependence Framework
Let $\epsilon^{*}_{0}$ and $\{\epsilon_{t}\}_{t\in\mathbb{Z}}$ be i.i.d. vectors in $\mathbb{R}^{b}$ and $\mathcal{F}_{t}=(\dots,\epsilon_{t-1},\epsilon_{t})$. We further let $\mathbf{x}_{l,t}=\left(x_{l,1t},\dots x_{l,pt}\right)^{T}$ be a $p$-dimensional process in condition $l$ where
$x_{l,jt}=R_{l,j}(\mathcal{F}_{t}).$
Note that $l$ indexes the condition, in this case 1 or 2, and $j$ indexes the variable, $1,\dots,p$. To measure the dependence of $\mathbf{x}_{l,t}$ on $\epsilon_{0}$ we can replace $\epsilon_{0}$ in $\mathcal{F}_{t}$ with an i.i.d. copy $\epsilon_{0}^{*}$. This gives $\mathcal{F}^{\prime}_{t}=(\dots,\epsilon_{-1},\epsilon_{0}^{*},\epsilon_{1},\dots,\epsilon_{t})$ and
$x^{\prime}_{l,jt}=R_{l,j}(\mathcal{F}^{\prime}_{t}).$
The dependence on $\epsilon_{0}$ can then be measured by
$\theta_{l,jt}=\left(\mathbb{E}\left|x_{l,jt}-x_{l,jt}^{\prime}\right|^{2}\right)^{1/2}$
where $\mathbb{E}\left|x\right|^{2}$ is the expected value of $\left|x\right|^{2}$. Controlling this dependence measure is essential for establishing the convergence rates of $\hat{f}_{l}$. The needed conditions on $\theta_{l,jt}$ are laid out in Assumption 2.
Assumption 2.
Assume for some constant $10<\rho<$ 1,
$\max_{l=\{1,2\}}\max_{j=1,\dots,p}\theta_{l,jt}=O(\rho^{t}),$
and for some constant $\kappa>0$, $C_{0}>0$,
$\max_{l=\{1,2\}}\max_{j=1,\dots,p}\mathbb{E}\left(\exp\left(\kappa|\mathbf{x}_{l,j0}|\right)\right)\leq C_{0}.$
Assumption 2 is essentially a restatement of Assumption 1 in Fiecas et al. (2019), except that we slightly generalize it to account for the fact that we have data from two separate conditions. This condition requires that the dependence on $\epsilon_{0}$ decreases geometrically as time increases. A similar condition is used in Wu and Zaffaroni (2018) and Liu and Wu (2010) and essentially requires that the dependence in the data on prior time points cannot be too strong. As noted in Shao and Wu (2007) and examples 1 and 2 of Liu and Wu (2010), many processes satisfy this constraint, including ARMA, ARMA-ARCH, and ARMA-GARCH processes as well as other nonlinear autoregressive processes.
Assumption 2 is satisfied if for example Assumption 1 from Fiecas et al. (2019) is satisfied for both conditions. For example if $\{\rho_{1},\kappa_{1},C_{0,1}\}$ and $\{\rho_{2},\kappa_{2},C_{0,2}\}$ are the parameters that satisfy Assumption 1 of Fiecas et al. (2019) for the data in conditions 1 and 2 respectively, then $\{\max(\rho_{1},\rho_{2}),\min(\kappa_{1},\kappa_{2}),\max(C_{0,1},C_{0,2})\}$ satisfy our Assumption 2. We are now ready to state the consistency of our estimator, $\hat{\Delta}$.
A.3 Proof of Theorem 1
Concentration Bound on $\hat{f}_{l}$ Around $f_{l}$.
We begin our proof by establishing a concentration inequality on
$\sup_{\{\lambda_{j}=2\pi j/n_{l},\lfloor(n_{l}-1)/2\rfloor\leq j\leq\lfloor n_{l}/2\rfloor\}}\|\hat{f}_{l}(\lambda_{j})-f_{l}(\lambda_{j})\|_{\infty}$
for each condition $i$ using Theorem 3.1 from (Fiecas et al., 2019). Under Assumption 2 we have that for any $\delta>0$ and $H>0$ and for $C_{1}>0$ depending only on $\kappa$ and $C_{0}$
$\mathbb{P}\left(\sup_{\{\lambda_{j}=2\pi j/n_{l},\lfloor(n_{l}-1)/2\rfloor\leq j\leq\lfloor n_{l}/2\rfloor\}}\|\hat{f}_{l}(\lambda_{j})-f_{l}(\lambda_{j})\|_{\infty}>\frac{C_{1}M_{n_{l}}}{n_{l}}+\frac{8n_{l}^{\delta}}{M_{n_{l}}^{1/2+\delta}}\right)\leq C_{2}p^{2}n_{l}^{-H}\,$
where $C_{2}$ is constant in $n_{l}$ but depends on $C_{1}$ and $H$. Note that this same inequality holds for both conditions $l=1,2$ because Assumption 2 holds for both conditions. We allow $n_{l}$ to differ between conditions for full generality. We can then select $M_{n_{l}}$ to minimize the deviations. That is we select $M_{n_{l}}$ to minimize $\frac{C_{1}M_{n_{l}}}{n_{l}}+\frac{8n_{l}^{\delta}}{M_{n_{l}}^{1/2+\delta}}$. The algebra is omitted but setting $\frac{\partial{}}{\partial{M_{n_{l}}}}\left[\frac{C_{1}M_{n_{l}}}{n_{l}}+\frac{8n_{l}^{\delta}}{M_{n_{l}}^{1/2+\delta}}\right]=0$ and solving we get $M^{*}_{n_{l}}=\left(\frac{8}{C_{1}}\left(\frac{1}{2}+\delta\right)\right)^{2/(3+2\delta)}n_{l}^{\frac{2+2\delta}{3+2\delta}}$. It can be shown that the second derivative is positive on $\delta>0$ so this is indeed a global minimum (on $\delta>0$). Plugging this back into the equation yields
$\frac{C_{1}M_{n_{l}}}{n_{l}}+\frac{8n_{l}^{\delta}}{M_{n_{l}}^{1/2+\delta}}\Bigg{|}_{M_{n_{l}}=M_{n_{l}}^{*}}=C_{1}^{\frac{1+2\delta}{3+2\delta}}8^{\frac{2}{3+2\delta}}\left(\left(\frac{1}{2}+\delta\right)^{\frac{2}{3+2\delta}}+\left(\frac{1}{2}+\delta\right)^{-\frac{1+2\delta}{3+2\delta}}\right)n_{l}^{-\frac{1}{3+2\delta}}\,.$
One can further show that $\left(\left(\frac{1}{2}+\delta\right)^{\frac{2}{3+2\delta}}+\left(\frac{1}{2}+\delta\right)^{-\frac{1+2\delta}{3+2\delta}}\right)$ is maximized at $\delta=1/2$ with a maximum value of $2$ (algebra omitted). Furthermore $C_{1}^{\frac{1+2\delta}{3+2\delta}}\leq C_{1}$ and $8^{\frac{2}{3+2\delta}}\leq 8^{2/3}$ for all $\delta>0$. Therefore we have that for all $\delta>0$,
$\frac{C_{1}M^{*}_{n_{l}}}{n_{l}}+\frac{8n_{l}^{\delta}}{(M^{*}_{n_{l}})^{1/2+\delta}}\leq C_{1}8^{2/3}2n_{l}^{-\frac{1}{3+2\delta}}\,.$
Letting $\delta\rightarrow 0$ the RHS of our concentration bound becomes $C_{1}8^{2/3}2n_{l}^{-\frac{1}{3}}$ and we get for any $H>0$ and for an optimal bandwidth choice of $M_{n_{l}}^{*}=O(n_{l}^{2/3})$ we get
$\mathbb{P}\left(\sup_{\{\lambda_{j}=2\pi j/n_{l},\lfloor(n_{l}-1)/2\rfloor\leq j\leq\lfloor n_{l}/2\rfloor\}}\|\hat{f}_{l}(\lambda_{j})-f_{l}(\lambda_{j})\|_{\infty}>C_{1}8^{2/3}2n_{l}^{-\frac{1}{3}}\right)\leq C_{2}p^{2}n_{l}^{-H}\,.$ (5)
Controlling Deviations of Sample Loss Around True Loss.
Throughout our theoretical analysis we will need to control the deviations of the second derivative of our sample loss around the true value. That is we will need to control
$\left\|0.5\left(\hat{\Sigma}_{1}\otimes\hat{\Sigma}_{2}+\hat{\Sigma}_{2}\otimes\hat{\Sigma}_{1}\right)-0.5\left(\Sigma_{1}\otimes\Sigma_{2}+\Sigma_{2}\otimes\Sigma_{1}\right)\right\|_{\infty}\,.$
Note that both
$\|\hat{\Sigma}_{2}\otimes\hat{\Sigma}_{1}-\Sigma_{2}\otimes\Sigma_{1}\|_{\infty},\|\hat{\Sigma}_{1}\otimes\hat{\Sigma}_{2}-\Sigma_{1}\otimes\Sigma_{2}\|_{\infty}\leq\|\Sigma_{1}\|_{\infty}\|\hat{\Sigma}_{2}-\Sigma_{2}\|_{\infty}+\|\Sigma_{2}\|_{\infty}\|\hat{\Sigma}_{1}-\Sigma_{1}\|_{\infty}+\|\hat{\Sigma}_{1}-\Sigma_{1}\|_{\infty}\|\hat{\Sigma}_{2}-\Sigma_{2}\|_{\infty}\,.$
First, note that $\|\Sigma_{l}\|_{\infty}\leq\|f_{l}\|_{\infty}$ and denote $C_{f}=\max(\|f_{1}\|_{\infty},\|f_{2}\|_{\infty})$. Using the optimal error bound on the smoothed periodograms from (5) we get that with probability $>1-2C_{2}p^{2}\min(n_{1},n_{2})^{-H}$,
$\|\hat{\Sigma}_{1}\otimes\hat{\Sigma}_{2}-\Sigma_{1}\otimes\Sigma_{2}\|_{\infty}$ $\leq C_{f}\left(C_{1}8^{3/2}2n_{1}^{-1/3}+C_{1}8^{3/2}2n_{2}^{-1/3}\right)+C_{1}^{2}8^{6/2}4n_{1}^{-1/3}n_{2}^{-1/3}$
$\leq C_{f}C_{1}8^{3/2}4\min(n_{1},n_{2})^{-1/3}+C_{1}^{2}8^{6/2}4\min(n_{1},n_{2})^{-2/3}\,.$
We can see that for $\min(n_{1},n_{2})\geq 8^{9/2}C_{1}^{3}/C_{f}^{3}$ the 2nd term on the RHS is smaller than the 1st term on the RHS so we can combine them to simplify the bound to be 2 times the first term on the RHS or
$\|\hat{\Sigma}_{1}\otimes\hat{\Sigma}_{2}-\Sigma_{1}\otimes\Sigma_{2}\|_{\infty}\leq\frac{C_{f}C_{1}8^{5/2}}{\min(n_{1},n_{2})^{1/3}}.$
This also holds for $\|\hat{\Sigma}_{2}\otimes\hat{\Sigma}_{1}-\Sigma_{2}\otimes\Sigma_{1}\|_{\infty}$. If we define
$\text{Event }\mathcal{A}:=\left\{\|\hat{\Sigma}_{1}\otimes\hat{\Sigma}_{2}-\Sigma_{1}\otimes\Sigma_{2}\|_{\infty}\leq\frac{C_{f}C_{1}8^{5/2}}{\min(n_{1},n_{2})^{1/3}}\right\}\cap\left\{\|\hat{\Sigma}_{2}\otimes\hat{\Sigma}_{1}-\Sigma_{2}\otimes\Sigma_{1}\|_{\infty}\leq\frac{C_{f}C_{1}8^{5/2}}{\min(n_{1},n_{2})^{1/3}}\right\}$
then from the above results, for $\min(n_{1},n_{2})\geq 8^{9/2}C_{1}^{3}/C_{f}^{3}$ Event $\mathcal{A}$ occurs with probability $>1-2C_{2}p^{2}\min(n_{1},n_{2})^{-H}$ and we have
$\left\|0.5\left(\hat{\Sigma}_{1}\otimes\hat{\Sigma}_{2}+\hat{\Sigma}_{2}\otimes\hat{\Sigma}_{1}\right)-0.5\left(\Sigma_{1}\otimes\Sigma_{2}+\Sigma_{2}\otimes\Sigma_{1}\right)\right\|_{\infty}\leq\frac{C_{f}C_{1}8^{5/2}}{\min(n_{1},n_{2})^{1/3}}\,.$ (6)
To prove the consistency of our estimator $\hat{\Delta}$ we use the framework of Negahban et al. (2012) which requires showing that our penalty, $\|\ \cdot\|_{1}$, is decomposable, that our D-trace loss function satisfies the restricted strong convexity (RSC) property, and that we suitably choose a penalty $\tau_{n_{1},n_{2}}$. Specifically we will use Theorem 1 from Negahban et al. (2012).
Decomposability of $\|\cdot\|_{1}$.
The decomposability of the penalty term is defined in Definition 1 of Negahban et al. (2012). For our loss, we let $S_{\Delta}$ be the support of $\Delta$, the true difference. Let ¯=$\overline{\mathcal{M}}=\left\{\theta:\mathbb{R}^{2p\times 2p}\big{|}\theta_{ij}\neq 0\text{ for all }(i,j)\in S_{\Delta}\right\}$ and ¯=$\overline{\mathcal{M}}^{\perp}=\left\{\gamma:\mathbb{R}^{2p\times 2p}\big{|}\gamma_{ij}=0\text{ for all }(i,j)\in S_{\Delta}\right\}$. Then $\|\cdot\|_{1}$ is decomposable since $\|\theta+\gamma\|_{1}=\|\theta\|_{1}+\|\gamma\|_{1}$ for all $\theta\in\overline{\mathcal{M}}$ and $\gamma\in\overline{\mathcal{M}}^{\perp}$.
RSC of D-trace Loss.
Next we must show that the D-trace loss satisfies the RSC condition from Definition 2 of Negahban et al. (2012). We must show that $\mathbf{m}^{T}\left(\nabla^{2}L(\Delta,\hat{\Sigma}_{1},\hat{\Sigma}_{2})\right)\mathbf{m}>\kappa_{L}\|\mathbf{m}\|_{2}^{2}$ for all $\mathbf{m}\in\mathcal{C}(S_{\Delta})$ where $\mathcal{C}(S_{\Delta})=\left\{\mathbf{m}\in\mathbb{R}^{4p^{2}}\big{|}\|\mathbf{m}_{S_{\Delta}^{c}}\|_{1}\leq 3\|\mathbf{m}_{S_{\Delta}}\|_{1}\right\}$ for some $\kappa_{L}$. The notation $\mathbf{m}_{S_{\Delta}}$ selects the entries of $\mathbf{m}$ that correspond to the non-zero indices of $\Delta$.
First note that for $\mathbf{m}\in\mathcal{C}(S_{\Delta})$, $\|\mathbf{m}\|_{1}\leq 4\|\mathbf{m}_{S_{\Delta}}\|_{1}\leq 4\sqrt{s_{\Delta}}\|\mathbf{m}_{S_{\Delta}}\|_{2}$ where $s_{\Delta}=|S_{\Delta}|$. Recall $\nabla^{2}L(\Delta,\hat{\Sigma}_{1},\hat{\Sigma}_{2})=0.5(\hat{\Sigma}_{1}\otimes\hat{\Sigma}_{2}+\hat{\Sigma_{2}}\otimes\hat{\Sigma}_{1})$. Showing RSC proceeds the same as in Wang et al. (2021).
$\begin{split}\mathbf{m}^{T}(\hat{\Sigma}_{1}\otimes\hat{\Sigma}_{2})\mathbf{m}&\geq\mathbf{m}^{T}(\Sigma_{1}\otimes\Sigma_{2})\mathbf{m}+\mathbf{m}^{T}\left(\hat{\Sigma}_{1}\otimes\hat{\Sigma}_{2}-\Sigma_{1}\otimes\Sigma_{2}\right)\mathbf{m}\\ &\geq\lambda_{\mathrm{min}}(\Sigma_{1})\lambda_{\mathrm{min}}(\Sigma_{2})\|\mathbf{m}\|_{2}^{2}-\left|\mathbf{m}^{T}\left(\hat{\Sigma}_{1}\otimes\hat{\Sigma}_{2}-\Sigma_{1}\otimes\Sigma_{2}\right)\mathbf{m}\right|\\ &\geq\lambda_{\mathrm{min}}(\Sigma_{1})\lambda_{\mathrm{min}}(\Sigma_{2})\|\mathbf{m}\|_{2}^{2}-\|\hat{\Sigma}_{1}\otimes\hat{\Sigma}_{2}-\Sigma_{1}\otimes\Sigma_{2}\|_{\infty}\|\mathbf{m}\|_{1}^{2}\\ &\geq\lambda_{\mathrm{min}}(\Sigma_{1})\lambda_{\mathrm{min}}(\Sigma_{2})\|\mathbf{m}\|_{2}^{2}-16s_{\Delta}\|\hat{\Sigma}_{1}\otimes\hat{\Sigma}_{2}-\Sigma_{1}\otimes\Sigma_{2}\|_{\infty}\|\mathbf{m}\|_{2}^{2}\,.\end{split}$ (7)
The last inequality follows from the fact that $\|\mathbf{m}\|_{1}=\|\mathbf{m}_{S_{\Delta}^{c}}\|_{1}+\|\mathbf{m}_{S_{\Delta}}\|_{1}\leq 4\|\mathbf{m}_{S_{\Delta}}\|_{1}\leq 4\sqrt{s_{\Delta}}\|\mathbf{m}_{S_{\Delta}}\|_{2}\leq 4\sqrt{s_{\Delta}}\|\mathbf{m}\|_{2}$. Similarly
$\mathbf{m}^{T}(\hat{\Sigma}_{2}\otimes\hat{\Sigma}_{1})\mathbf{m}\geq\lambda_{\mathrm{min}}(\Sigma_{1})\lambda_{\mathrm{min}}(\Sigma_{2})\|\mathbf{m}\|_{2}^{2}-16s_{\Delta}\|\hat{\Sigma}_{2}\otimes\hat{\Sigma}_{1}-\Sigma_{2}\otimes\Sigma_{1}\|_{\infty}\|\mathbf{m}\|_{2}^{2}\,.$
Recall we showed that for $\min(n_{1},n_{2})\geq 8^{9/2}C_{1}^{3}/C_{f}^{3}$, Event $\mathcal{A}$ occurs with probability $>1-2C_{2}p^{2}\min(n_{1},n_{2})^{-H}$. Using this and the fact that $\lambda_{\mathrm{min}}(\Sigma_{l})=\lambda_{\mathrm{min}}(f_{l})$ we have that
$\mathbf{m}^{T}0.5(\hat{\Sigma}_{1}\otimes\hat{\Sigma}_{2}+\hat{\Sigma}_{2}\otimes\hat{\Sigma}_{1})\mathbf{m}\geq\lambda_{\mathrm{min}}(f_{1})\lambda_{\mathrm{min}}(f_{2})\|\mathbf{m}\|_{2}^{2}-16s_{\Delta}\frac{C_{f}C_{1}8^{5/2}}{\min(n_{1},n_{2})^{1/3}}\|\mathbf{m}\|_{2}^{2}$
So for $\min(n_{1},n_{2})\geq\left(\frac{32s_{\Delta}C_{f}C_{1}8^{5/2}}{\lambda_{\mathrm{min}}(f_{1})\lambda_{\mathrm{min}}(f_{2})}\right)^{3}$ then the 2nd term on the RHS is $\leq$ 0.5*(1st term on RHS). Thus we get
$\mathbf{m}^{T}0.5(\hat{\Sigma}_{1}\otimes\hat{\Sigma}_{2}+\hat{\Sigma}_{2}\otimes\hat{\Sigma}_{1})\mathbf{m}\geq\frac{\lambda_{\mathrm{min}}(f_{1})\lambda_{\mathrm{min}}(f_{2})}{2}\|\mathbf{m}\|_{2}^{2}$
Therefore we conclude that for $\min(n_{1},n_{2})\geq\max\left(8^{9/2}C_{1}^{3}/C_{f}^{3},\left(\frac{32s_{\Delta}C_{f}C_{1}8^{5/2}}{\lambda_{\mathrm{min}}(f_{1})\lambda_{\mathrm{min}}(f_{2})}\right)^{3}\right)$ with probability ¿ $1-2C_{2}p^{2}\min(n_{1},n_{2})^{-H}$ RSC holds with $\kappa_{L}=\lambda_{\mathrm{min}}(f_{1})\lambda_{\mathrm{min}}(f_{2})/2$.
Choice of Penalty $\boldsymbol{\tau_{n_{1},n_{2}}}$.
The last ingredient in our theoretical analysis is the choice of penalty scale $\tau_{n_{1},n_{2}}$. Theorem 1 in Negahban et al. (2012) we need $\tau_{n_{1},n_{2}}\geq 2\|\nabla L_{D}(\Delta,\hat{\Sigma}_{1},\hat{\Sigma}_{2})\|_{\infty}$. We know that $2\|\nabla L_{D}(\Delta,\hat{\Sigma}_{1},\hat{\Sigma}_{2})\|_{\infty}=2\|0.5\left(\hat{\Sigma}_{1}\Delta\hat{\Sigma}_{2}+\hat{\Sigma}_{2}\Delta\hat{\Sigma}_{1}\right)-\left(\hat{\Sigma}_{1}-\hat{\Sigma}_{2}\right)\|_{\infty}.$
Let $\Gamma=0.5\left(\Sigma_{2}\otimes\Sigma_{1}+\Sigma_{1}\otimes\Sigma_{2}\right)$ and $\hat{\Gamma}=0.5\left(\hat{\Sigma}_{2}\otimes\hat{\Sigma}_{1}+\hat{\Sigma}_{1}\otimes\hat{\Sigma}_{2}\right).$
Then we write
$\|0.5\left(\hat{\Sigma}_{1}\Delta\hat{\Sigma}_{2}+\hat{\Sigma}_{2}\Delta\hat{\Sigma}_{1}\right)-\left(\hat{\Sigma}_{1}-\hat{\Sigma}_{2}\right)\|_{\infty}$ $=\|\hat{\Gamma}\text{vec}(\Delta)-\left(\text{vec}(\hat{\Sigma}_{1})-\text{vec}(\hat{\Sigma}_{2})\right)\|_{\infty}$
$=\|(\hat{\Gamma}-\Gamma)\text{vec}(\Delta)-\left(\text{vec}(\hat{\Sigma}_{1})-\text{vec}(\Sigma_{1})\right)-\left(\text{vec}(\hat{\Sigma}_{2})-\text{vec}(\Sigma_{2})\right)\|_{\infty}$
$\leq\|(\hat{\Gamma}-\Gamma)\text{vec}(\Delta)\|_{\infty}+\|\hat{\Sigma}_{1}-\Sigma_{1}\|_{\infty}+\|\hat{\Sigma}_{2}-\Sigma_{2}\|_{\infty}$
$\leq\|\hat{\Gamma}-\Gamma\|_{\infty}\|\Delta\|_{1}+\|\hat{\Sigma}_{1}-\Sigma_{1}\|_{\infty}+\|\hat{\Sigma}_{2}-\Sigma_{2}\|_{\infty}\,.$
Note that we showed in (6) that for $\min(n_{1},n_{2})\geq\frac{C_{1}^{3}}{C_{f}^{3}}8^{9/2}$ we have with probability $>1-2C_{2}p^{2}\min(n_{1},n_{2})^{-H}$,
$\|\hat{\Gamma}-\Gamma\|_{\infty}\leq\frac{C_{1}C_{f}8^{5/2}}{\min(n_{1},n_{2})^{1/3}}\,.$
Thus we get
$\left\|0.5\left(\hat{\Sigma}_{1}\Delta\hat{\Sigma}_{2}+\hat{\Sigma}_{2}\Delta\hat{\Sigma}_{1}\right)-\left(\hat{\Sigma}_{1}-\hat{\Sigma}_{2}\right)\right\|_{\infty}$ $\leq\frac{C_{1}C_{f}8^{5/2}}{\min(n_{1},n_{2})^{1/3}}\|\Delta\|_{1}+\frac{C_{1}8^{3/2}2}{n_{1}^{1/3}}+\frac{C_{1}8^{3/2}2}{n_{2}^{1/3}}$
$\leq\frac{\left(C_{1}C_{f}8^{5/2}\|\Delta\|_{1}+4C_{1}8^{3/2}\right)}{\min(n_{1},n_{2})^{1/3}}\,.$
Therefore if we choose
$\tau_{n_{1},n_{2}}\geq\frac{2\left(C_{1}C_{f}8^{5/2}\|\Delta\|_{1}+4C_{1}8^{3/2}\right)}{\min(n_{1},n_{2})^{1/3}}$
we will satisfy $\tau_{n_{1},n_{2}}\geq 2\|\nabla L_{D}(\Delta,\hat{\Sigma}_{1},\hat{\Sigma}_{2})\|_{\infty}$ with probability $>1-2C_{2}p^{2}\min(n_{1},n_{2})^{-H}$ when $\min(n_{1},n_{2})\geq\frac{C_{1}^{3}}{C_{f}^{3}}8^{9/2}$.
Convergence Rate for Difference in Networks.
Now we are ready to state the result which uses Corollary 1 of Negahban et al. (2012).
Under Assumption 2 for $\tau_{n_{1},n_{2}}\geq\frac{2}{\min(n_{1},n_{2})^{1/3}}\left(C_{1}C_{f}8^{5/2}\|\Delta\|_{1}+4C_{1}8^{3/2}\right)$,
$\min(n_{1},n_{2})\geq\max\left(8^{9/2}C_{1}^{3}/C_{f}^{3},\left(\frac{32s_{\Delta}C_{f}C_{1}8^{5/2}}{\lambda_{\mathrm{min}}(f_{1})\lambda_{\mathrm{min}}(f_{2})}\right)^{3}\right)$, and $M_{n,l}=O(n_{l}^{2/3})$ we have with probability greater than $1-C_{2}p^{2}\min(n_{1},n_{2})^{-H}$ for any $H>0$,
$\|\hat{\Delta}-\Delta\|_{F}\leq\frac{\sqrt{18s_{\Delta}}\tau_{n_{1},n_{2}}}{\lambda_{\mathrm{min}}(f_{1})\lambda_{\mathrm{min}}(f_{2})}$
Since $\tau_{n_{1},n_{2}}$ is of order $\min(n_{1},n_{2})^{-1/3}$, convergence of our D-trace estimator is also of order $\min(n_{1},n_{2})^{-1/3}$.
A.4 Simulation Specifics and Results
We generate data in condition $l$ as
$\mathbf{x}_{l,t}=A_{l}\mathbf{x}_{l,t-1}+\boldsymbol{\epsilon}_{l,t}\,$
where $\boldsymbol{\epsilon}_{l,t}\sim N_{p}\left(\mathbf{0},I_{p}\right)$. From Sun et al. (2018), the spectral density at frequency $\lambda$ is known to be
$f_{l}(\lambda)=\frac{1}{2\pi}\left(\mathcal{A}_{l}(e^{-i\lambda})\right)^{-1}I_{p}\left(\left(\mathcal{A}_{l}(e^{-i\lambda})\right)^{-1}\right)^{H},$
where $\mathcal{A}_{l}(z)=I_{p}-A_{l}z$. When the transition matrix $A_{l}$ is block diagonal, the spectral density $f_{l}(\lambda)$ is also block diagonal. Since the inverse of a block diagonal matrix can be computed block by block we can easily generate a sparse difference matrix by enforcing the transition matrix in conditions $1$ and $2$ to be the same except for a small block. For example, with $p=54$ if we generate $A_{1}$ and $A_{2}$ as block diagonal with the same $51\times 51$ block and only differing in the final $3\times 3$ block, then their spectral densities will only differ in this last $3\times 3$ block. When converting to the real space, the expanded $\Sigma_{i}\in(2p\times 2p)$ and thus the true difference $\Delta=\Sigma_{1}-\Sigma_{2}$ will differ in four $3\times 3$ blocks.
For simulation setting 1, the transition matrix is the same as in (Sun et al., 2018). Specifically, it consists of $18$ blocks of dimension $3\times 3$, where each block is $\begin{bmatrix}0.5&0.9&0\\ 0&0.5&0.9\\ 0&0&0.5\end{bmatrix}$. In both simulation setting 2 and setting 3, 60% of the coefficients in the $3\times 3$ block were randomly drawn from either a $\mathrm{Uniform}(-0.5,-0.2)$ or a $\mathrm{Uniform}(0.2,0.5)$ each with equal probability. In the second and third setting, 40% and 5% respectively of the entries of the larger block were randomly drawn from a $\mathrm{Uniform}(-0.5,-0.2)$ and $\mathrm{Uniform}(0.2,0.5)$ each with equal probability. This corresponded to 60% and 95% sparsity respectively. The number of non-zero entries in the difference in expanded inverse spectral densities, which we will refer to as edges, varies by frequency but is almost always 22 for Simulation 1, 14 for Simulation 2, and 28 for Simulation 3.
For all settings $100$ evenly spaced Fourier frequencies from 00 to $\pi-n^{-1}$ was used. These were used as the spectral density is conjugate symmetric around 00. In the case of $n=100$ only $50$ Fourier frequencies were used as there are only 50 Fourier frequencies from 00 to $\pi-n^{-1}$. Given the availability of results from multiple frequencies, each simulation was run once and the results are averaged across all frequencies. The standard error of each metric was also computed across frequencies and are reported in parentheses.
To solve the SDD estimation problem (1), the alternating direction method of multipliers (ADMM) algorithm from Yuan et al. (2017) was used. Specifically, we used the L1_dts function in the Difdtl package available on GitHub at SusanYuan/Difdtl (GPL ($\geq$ 2)). To generate the sequence of penalties, $\{\tau_{n_{1},n_{2}}\}$, we used 20 values on a log-linear scale from $0.001*\tau_{n_{1},n_{2},\mathrm{max}}$ to $\tau_{n_{1},n_{2},\mathrm{max}}$ where $\tau_{n_{1},n_{2},\mathrm{max}}$ represents the minimum value where all entries of $\hat{\Delta}$ are 0. In this case, $\tau_{n_{1},n_{2},\mathrm{max}}=2\max(|\hat{\Sigma}_{1}-\hat{\Sigma}_{2}|)$. The penalty $\tau_{n_{1},n_{2}}^{*}$ was chosen as the $\tau_{n_{1},n_{2}}$ that minimizes the BIC which is computed as
$\mathrm{BIC}(\hat{\Delta})=2n\left\|\frac{1}{2}\left(\hat{\Sigma}_{1}\hat{\Delta}\hat{\Sigma}_{2}+\hat{\Sigma}_{2}\hat{\Delta}\hat{\Sigma}_{1}-\hat{\Sigma}_{2}+\hat{\Sigma}_{1}\right)\right\|_{F}+\log(2n)\|\hat{\Delta}\|_{0}\,.$
Let TP, FN, TN, FP denote the true positive, false negative, true negative, and false positive edges identified by either SDD or the Naïve method, respectively. A value greater than $1\times 10^{-6}$ in absolute value was considered an edge. The metrics are defined as follows
$\mathrm{Precision}=\frac{\mathrm{TP}}{\mathrm{TP}+\mathrm{FP}}$ $\mathrm{Recall}=\frac{\mathrm{TP}}{\mathrm{TP}+\mathrm{FN}}$
$\mathrm{Accuracy}=\frac{\mathrm{TP}+\mathrm{TN}}{4p^{2}}$ $\mathrm{RRMSE}=\sqrt{\frac{\sum_{i,j}(\hat{\Delta}_{i,j}-\Delta_{i,j})^{2}}{\sum_{i,j}(\Delta_{i,j})^{2}}},$
where $\hat{\Delta}$ represents the difference estimator, either SDD or the Naïve difference, $\Delta$ is the true difference. The denominator of the accuracy measure is $4p^{2}$ as expanding a $p\times p$ spectral density to the real space gives a $2p\times 2p$ matrix which has $4p^{2}$ entries. We also report the number of average number of true edges and the average number of estimated edges across frequencies.
Figure 8: Simulation 1 Precision and Recall. Results are reported as mean (dots) and SE (vertical lines) where the mean and SE are taken across all frequencies for a given sample size $n$.
Figure 9: Simulation 2 Precision and Recall. Results are reported as mean (dots) and SE (vertical lines) where the mean and SE are taken across all frequencies for a given sample size $n$.
Figure 10: Simulation 3 Accuracy and RRMSE. Results are reported as mean (dots) and SE (vertical lines) where the mean and SE are taken across all frequencies for a given sample size $n$.
Figure 11: Simulation 3 Precision and Recall. Results are reported as mean (dots) and SE (vertical lines) where the mean and SE are taken across all frequencies for a given sample size $n$.
Table 1: Simulation 1. Results are reported as Mean (SE) where the mean and SE are taken across all frequencies for a given sample size $n$.
SDD
n # True edges # Est edges Precision Recall Accuracy RRMSE
100 21.6 (2.22) 62.8 (71.0) 0.42 (0.47) 0.16 (0.14) 0.99 (0.01) 0.96 (0.04)
200 21.8 (1.59) 134 (183) 0.58 (0.48) 0.37 (0.17) 0.99 (0.02) 0.87 (0.08)
500 21.9 (0.79) 306 (327) 0.39 (0.44) 0.68 (0.25) 0.97 (0.03) 0.68 (0.17)
1000 21.9 (0.79) 447 (511) 0.40 (0.44) 0.78 (0.25) 0.96 (0.04) 0.49 (0.17)
2000 21.9 (0.79) 647 (742) 0.35 (0.42) 0.85 (0.22) 0.95 (0.06) 0.42 (0.12)
Naïve
659 (323) 0.03 (0.02) 0.62 (0.15) 0.94 (0.03) 0.95 (0.04)
1077 (566) 0.03 (0.02) 0.79 (0.19) 0.91 (0.05) 0.90 (0.05)
1775 (718) 0.01 (0.01) 0.91 (0.13) 0.85 (0.06) 0.80 (0.07)
2289 (1035) 0.01 (0.01) 0.99 (0.05) 0.81 (0.09) 0.69 (0.10)
3158 (1500) 0.01 (0.00) 0.99 (0.04) 0.73 (0.13) 0.58 (0.08)
Hard threshold
0.82 (1.42) 0.39 (0.48) 0.02 (0.04) 1.00 (0.00) 1.59 (0.63)
2.06 (5.58) 0.66 (0.45) 0.06 (0.10) 1.00 (0.00) 1.01 (0.08)
7.59 (9.74) 0.77 (0.40) 0.31 (0.37) 1.00 (0.00) 0.78 (0.24)
85.7 (559) 0.82 (0.38) 0.40 (0.38) 0.99 (0.05) 0.64 (0.42)
Table 2: Simulation 2. Results are reported as Mean (SE) where the mean and SE are taken across all frequencies for a given sample size $n$.
SDD
n # True edges # Est edges Precision Recall Accuracy RRMSE
100 13.7 (1.48) 3.32 (4.66) 0.00 (0.00) 0.00 (0.00) 1.00 (0.00) 1.00 (0.00)
200 13.9 (1.06) 2.62 (6.27) 0.00 (0.00) 0.00 (0.00) 1.00 (0.00) 1.00 (0.00)
500 14.0 (0.40) 18.2 (15.8) 0.01 (0.05) 0.02 (0.07) 1.00 (0.00) 0.99 (0.04)
1000 14.0 (0.40) 23.5 (39.5) 0.29 (0.43) 0.13 (0.12) 1.00 (0.00) 0.97 (0.06)
2000 14.0 (0.40) 113 (73.7) 0.08 (0.14) 0.38 (0.13) 0.99 (0.01) 0.71 (0.11)
Naïve
259 (76.2) 0.02 (0.01) 0.42 (0.06) 0.98 (0.01) 1.02 (0.04)
609 (460) 0.02 (0.02) 0.43 (0.05) 0.95 (0.04) 1.01 (0.07)
3566 (560) 0.00 (0.00) 0.54 (0.14) 0.69 (0.05) 1.20 (0.27)
6487 (579) 0.00 (0.00) 0.79 (0.21) 0.44 (0.05) 1.58 (0.44)
9145 (977) 0.00 (0.00) 0.95 (0.11) 0.22 (0.08) 2.29 (0.38)
Hard threshold
0.72 (0.51) 0.06 (0.24) 0.00 (0.02) 1.00 (0.00) 3.05 (1.79)
0.93 (0.80) 0.46 (0.50) 0.04 (0.06) 1.00 (0.00) 1.09 (0.20)
1.05 (0.82) 0.78 (0.41) 0.07 (0.06) 1.00 (0.00) 0.93 (0.10)
1.30 (0.97) 0.85 (0.36) 0.09 (0.07) 1.00 (0.00) 0.87 (0.11)
Table 3: Simulation 3. Results are reported as Mean (SE) where the mean and SE are taken across all frequencies for a given sample size $n$.
SDD
n # True edges # Est edges Precision Recall Accuracy RRMSE
100 27.4 (2.74) 1.76 (0.85) 0.82 (0.38) 0.06 (0.03) 1.00 (0.00) 0.99 (0.02)
200 27.7 (1.97) 2.30 (1.36) 0.83 (0.37) 0.08 (0.05) 1.00 (0.00) 0.97 (0.03)
500 27.9 (1.19) 20.6 (29.6) 0.55 (0.40) 0.19 (0.17) 1.00 (0.00) 0.87 (0.14)
1000 27.9 (1.19) 13.8 (20.0) 0.79 (0.33) 0.28 (0.22) 1.00 (0.00) 0.79 (0.19)
2000 27.9 (1.19) 61.9 (55.8) 0.57 (0.29) 0.78 (0.22) 1.00 (0.00) 0.52 (0.13)
Naïve
308 (89.2) 0.02 (0.01) 0.26 (0.10) 0.97 (0.01) 0.94 (0.04)
563 (201) 0.02 (0.01) 0.36 (0.14) 0.95 (0.02) 0.91 (0.06)
2331 (344) 0.01 (0.00) 0.88 (0.15) 0.80 (0.03) 0.81 (0.07)
3532 (496) 0.01 (0.00) 0.93 (0.13) 0.70 (0.04) 0.71 (0.08)
5030 (615) 0.01 (0.00) 0.97 (0.08) 0.57 (0.05) 0.61 (0.07)
Hard threshold
0.58 (0.53) 0.29 (0.45) 0.01 (0.02) 1.00 (0.00) 1.74 (0.90)
1.01 (1.75) 0.65 (0.47) 0.04 (0.06) 1.00 (0.00) 1.01 (0.04)
0.96 (1.02) 0.66 (0.47) 0.03 (0.04) 1.00 (0.00) 0.96 (0.06)
7.33 (9.40) 0.83 (0.37) 0.27 (0.34) 1.00 (0.00) 0.69 (0.34)
Table 4: Simulation 1. Difference Between SDD and Naïve Method for Each Metric.
n # Est edges Precision Recall Accuracy RRMSE
100 -597 (265) 0.39 (0.45) -0.46 (0.25) 0.05 (0.02) 0.01 (0.06)
200 -943 (431) 0.55 (0.47) -0.42 (0.20) 0.08 (0.04) -0.03 (0.07)
500 -1469 (606) 0.38 (0.44) -0.23 (0.21) 0.13 (0.05) -0.12 (0.18)
1000 -1842 (808) 0.39 (0.44) -0.21 (0.26) 0.16 (0.07) -0.20 (0.21)
2000 -2511 (1012) 0.35 (0.42) -0.14 (0.22) 0.21 (0.09) -0.16 (0.17)
Table 5: Simulation 1. Difference Between SDD and Hard Thresholding Method for Each Metric.
n # Est edges Precision Recall Accuracy RRMSE
100 - - - - -
200 133 (183) 0.19 (0.79) 0.35 (0.18) -0.01 (0.02) -0.72 (0.66)
500 304 (328) -0.26 (0.64) 0.61 (0.26) -0.02 (0.03) -0.33 (0.20)
1000 439 (514) -0.36 (0.61) 0.47 (0.37) -0.04 (0.04) -0.29 (0.28)
2000 561 (852) -0.46 (0.53) 0.45 (0.40) -0.05 (0.07) -0.22 (0.44)
Table 6: Simulation 2. Difference Between SDD and Naïve Method for Each Metric.
n # Est edges Precision Recall Accuracy RRMSE
100 -255 (73.4) -0.02 (0.01) -0.42 (0.06) 0.02 (0.01) -0.02 (0.04)
200 -607 (457) -0.02 (0.02) -0.43 (0.05) 0.05 (0.04) -0.01 (0.07)
500 -3548 (562) 0.01 (0.05) -0.52 (0.15) 0.30 (0.05) -0.21 (0.27)
1000 -6463 (585) 0.28 (0.43) -0.66 (0.23) 0.55 (0.05) -0.61 (0.44)
2000 -9032 (1021) 0.08 (0.14) -0.58 (0.14) 0.77 (0.09) -1.57 (0.37)
Table 7: Simulation 2. Difference Between SDD and Hard Thresholding Method for Each Metric.
n # Est edges Precision Recall Accuracy RRMSE
100 - - - - -
200 1.90 (6.20) -0.06 (0.24) -0.00 (0.02) -0.00 (0.00) -2.05 (1.79)
500 17.3 (15.9) -0.45 (0.49) -0.02 (0.08) -0.00 (0.00) -0.10 (0.20)
1000 22.5 (39.3) -0.49 (0.58) 0.06 (0.13) -0.00 (0.00) 0.05 (0.11)
2000 112 (73.7) -0.77 (0.38) 0.28 (0.14) -0.01 (0.01) -0.16 (0.15)
Table 8: Simulation 3. Difference Between SDD and Naïve Method for Each Metric.
n # Est edges Precision Recall Accuracy RRMSE
100 -306 (89.1) 0.80 (0.38) -0.20 (0.09) 0.03 (0.01) 0.04 (0.03)
200 -561 (200) 0.81 (0.37) -0.27 (0.12) 0.05 (0.02) 0.06 (0.04)
500 -2310 (336) 0.54 (0.40) -0.69 (0.22) 0.19 (0.03) 0.06 (0.10)
1000 -3518 (496) 0.78 (0.33) -0.65 (0.26) 0.30 (0.04) 0.08 (0.13)
2000 -4968 (618) 0.56 (0.29) -0.19 (0.20) 0.42 (0.05) -0.09 (0.11)
Table 9: Simulation 3. Difference Between SDD and Hard Thresholding Method for Each Metric.
n # Est edges Precision Recall Accuracy RRMSE
100 - - - - -
200 1.72 (1.50) 0.54 (0.50) 0.07 (0.05) 0.00 (0.00) -0.77 (0.91)
500 19.6 (29.7) -0.10 (0.59) 0.16 (0.18) -0.00 (0.00) -0.14 (0.14)
1000 12.9 (20.1) 0.13 (0.60) 0.25 (0.22) 0.00 (0.00) -0.17 (0.19)
2000 54.5 (58.5) -0.26 (0.44) 0.51 (0.37) -0.00 (0.00) -0.17 (0.33)
A.5 Additional $\mu\mathrm{ECoG}$ Analyses
Figure 12: Relationship Between Probability of Edge Change and Distance Between Electrodes Among Control Sessions. Lines are generated by smoothing using a GAM. Shaded areas represent 95% confidence intervals. Electrodes that are close together have a higher probability of edge change.
Figure 13: Relationship Between Average Distance of Electrode Pair to Closest Stimulation Site and Distance Between Electrodes Among All Sessions. Lines are generated by smoothing using a LOESS smoother. Shaded areas represent 95% confidence intervals. Electrodes pairs that are either very close or very far from stimulation sites are close together.
Figure 13 shows that electrode pairs that have a small or large mean stim distance also have small distance between them. This makes sense as if two electrodes are close together both are either far or close to the stim sites. Taken with Figure 12, this means that even in control sessions we expect a U-shaped relationship between distance to stimulation site and probability of edge change.